### v1

# DarkAI: Reconstructing the density, velocity and tidal field of dark matter from DESI-like bright galaxy sample

Abstract
Reconstructing the dark matter density, velocity, and tidal (MTV) fields from galaxy surveys is essential for advancing our understanding of the large-scale structure of the Universe. In this work, we present a machine learning-based framework using a UNet convolutional neural network to reconstruct the MTV fields from mock samples of the DESI bright galaxy survey within the redshift range $0.40.1<z<$ 0.4. Our approach accounts for realistic observational effects, including geometric selection, flux-limited data, and redshift space distortion (RSD) effects, thereby improving the fidelity of the reconstructed fields. Testing on mock galaxy catalogs generated from the Jiutian N-body simulation, our method achieves significant accuracy level. The reconstructed density field exhibits strong consistency with the true field, effectively eliminating most RSD effects and achieving a cross-correlation power spectrum coefficient greater than 0.985 on scales with $k<0.1\,h\,\mathrm{Mpc}^{-1}$. The velocity field reconstruction accurately captures large-scale coherent flows and small-scale turbulent features, exhibiting slopes of the grid-to-grid relationship close to unity and a scatter below $\sim$100 $\mathrm{km}\,\mathrm{s}^{-1}$. Additionally, the tidal field is reconstructed without bias, successfully recovering the features of the large-scale cosmic web, including clusters, filaments, sheets, and voids. Our results confirm that the proposed framework effectively captures the large-scale distribution and dynamics of dark matter while addressing key systematic challenges. These advancements provide a reliable and robust tool for analyzing current and future galaxy surveys, paving the way for new insights into cosmic structure formation and evolution.
Large-scale structure of the universe(902) — Dark matter distribution(356) — Redshift surveys(1378)—Convolutional neural networks(1938)
software: Astropy (Astropy Collaboration et al., 2018, 2022), NumPy (van der Walt et al., 2011; Harris et al., 2020), Matplotlib (Hunter, 2007), Tensorflow(Abadi et al., 2015), Pylians(Villaescusa-Navarro, 2018)
1 Introduction
Understanding the distribution and dynamics of matter in the Universe provides critical insights into gravitational interactions and cosmic flows. This knowledge is essential for exploring the processes that drive the formation and evolution of cosmic structures and galaxies. To achieve this, galaxy redshift surveys, such as the 2-degree Field Galaxy Redshift Survey (Colless et al., 2001) and the Sloan Digital Sky Survey (SDSS) (York et al., 2000), offer an opportunity to probe the large-scale mass distribution of the Universe. In the coming decade, next-generation galaxy surveys, including the Dark Energy Spectroscopic Instrument (DESI) (DESI Collaboration et al., 2016a, b), the Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST) (Ivezić et al., 2019), EUCLID (Laureijs et al., 2011), the Nancy Grace Roman Space Telescope (Akeson et al., 2019), and the Chinese Space Station Telescope (CSST) (Zhan, 2011; Gong et al., 2019), are poised to map an unprecedented volume of the Universe with different redshift measurement accuracy. These surveys will enable transformative advances in our understanding of cosmic structure and evolution, emphasizing the importance of developing robust and efficient methods to reconstruct the three-dimensional dark-matter density, velocity, and tidal (MTV) field from the vast galaxy survey datasets.
Reconstructing the MTV field from observational data is essential for addressing key scientific questions. Constrained simulations, such as those in the Local Universe (e.g., Wang et al., 2016; Tweed et al., 2017), require accurate initial conditions derived from the density field to reproduce observed structures. Local non-Gaussianity in the cosmic web can be quantified through the MTV field, offering insights into primordial fluctuations (Desjacques & Seljak, 2010; Smith et al., 2018; Münchmeyer et al., 2019). Filament spin studies, such as alignment of the galaxy angular momentum with filaments (e.g., Tempel & Libeskind, 2013; Wang et al., 2021), rely on the reconstructed tidal field for precise measurements. Filament lensing studies, mapping dark matter in filaments (e.g., HyeongHan et al., 2024), are similarly dependent on accurate reconstructions. Moreover, correcting supernova distance measurements for mass density and peculiar velocities (Cooray & Caldwell, 2006; Hui & Greene, 2006; Davis et al., 2011; Huang et al., 2024) and mapping baryon distributions using the velocity field (e.g., Dolag & Sunyaev, 2013; Yu & Zhu, 2019) underscore the broad utility of MTV reconstructions.
Reconstructing the cosmic density and velocity fields has been conducted earlier from several redshift surveys (Fisher et al., 1995; Zaroubi et al., 1995; Schmoldt et al., 1999; Mathis et al., 2002; Erdoǧdu et al., 2004; Wang et al., 2009, 2013). Early efforts (Fisher et al., 1995; Zaroubi et al., 1995) relied on smoothing and normalizing galaxy data to approximate large-scale density fields, with methods such as Wiener reconstruction minimizing errors through linear combinations of galaxy density values. Other approaches (Wang et al., 2009, 2012, 2013) leveraged dark-matter halos or galaxy groups, assuming a linear bias between halo and mass density fields. However, these methods are quite complicated. The assumption of a scale-independent bias parameter lacks robust physical justification and is often motivated by simplicity, reducing its reliability in capturing complex relationships (e.g., Coles & Erdogdu, 2007; Simon & Hilbert, 2018). Moreover, linear theory assumptions, which posit a simple proportionality between gravity and velocity fields, fail in the nonlinear clustering regime due to phenomena like shell crossing, where the one-to-one correspondence between density and velocity fields breaks down. Additionally, redshift space distortions (RSDs) (Sargent & Turner, 1977; Davis & Peebles, 1983; Kaiser, 1987; Hamilton, 1992), influenced by the Kaiser effect on large scales (Kaiser, 1987) and the Finger-of-God effect on small scales (Jackson, 1972; Tully & Fisher, 1978), intertwine complex velocity and density field distortions, further challenging accurate reconstructions. Methods based on linear theory or the Zeldovich approximation are particularly limited in high-density regions, emphasizing the need for innovative strategies to fully leverage redshift survey data in mapping the Universe’s large-scale structure.
Several studies have employed machine learning (ML) techniques to reconstruct dark matter density and velocity fields, demonstrating significant potential for advancing the field. A UNet-based neural network was used to reconstruct peculiar velocity fields from redshift-space distributions of dark matter halos, achieving high accuracy at multiple scales(Wu et al., 2021, 2023; Qin et al., 2023; Wang et al., 2024; Xiao et al., 2024). However, these approaches face notable limitations. First, these approaches often operate within idealized simulation environments, with limitations such as fixed box volumes, periodic boundary conditions, and the exclusion of observational systematics like geometric selection and flux-limited effects. Second, reconstruction based on the halo distributions compared to the galaxies requires additional bias modeling to bridge the gap between galaxy and dark matter. Addressing these limitations is essential for extending the ML-based reconstruction to real-world cosmological observations.
In this paper, we aim to develop a robust ML-based framework to reconstruct the MTV fields from DESI-like bright galaxy samples. In Wang et al. (2024, hereafter Wang24), the first paper in this series, we present a deep-learning technique for reconstructing the dark-matter density field from the redshift-space distribution of dark-matter halos. The reconstruction was clearly proved to be able to correct for redshift-space distortions and is unaffected by the different cosmologies between the training and test samples. It is also tested that the application of the UNet-reconstructed density field can accurately infer the velocity & tidal field accurately, providing better results compared to the traditional approach based on the linear bias model. Here we improve upon the reconstruction method of Wang24 by addressing realistic observation challenges, including accounting for geometric selection, varying redshift distributions, and flux-limited observational biases. By incorporating systematic effects into the ML model, we aim to enhance the fidelity of the reconstructed fields and expand their applicability to real survey data. This study concentrates solely on reconstruction using the distribution of galaxies. We will address reconstruction involving groups or halos derived from observational data in a future probe.
The structure of this paper is as follows. Section 2 provides an overview of the data and mock galaxy catalogs used for training and testing. Section 3 describes the ML-based reconstruction framework, including network architecture and training protocols. Section 4 evaluates the performance of reconstruction on mock DESI-like catalogs, addressing systematic effects and observational biases. Finally, Section 5 presents a summary and prospects for future work.
2 data
To support the DESI galaxy survey, we generate mock galaxy samples along with their corresponding dark matter density and velocity fields within a DESI-like light cone. These mock data sets are essential for studying the mapping between galaxies and the underlying large-scale structure of the universe, as well as for testing and validating cosmological analysis pipelines. The mock samples are constructed through a multi-step process that includes N-body simulations, halo and subhalo identification, galaxy population modeling, and the incorporation of survey selection effects. This section outlines the methodology utilized to generate these mock samples, based on the procedures described in Gu et al. (2024).
The foundation of our mock galaxy samples is the Jiutian N-body simulation, which was run at the High-Performance Computing Center at Kunshan using L-GADGET, a memory-optimized version of the GADGET2 code (Springel, 2005). The Jiutian simulation describes the distribution of $6144^{3}$ dark-matter particles in a periodic box of $1000\>h^{-1}{\rm{Mpc}}$, with a particle mass of $3.723\times 10^{8}h^{-1}M_{\odot}$. The cosmological parameters were set to be consistent with the Planck2018 results (Planck Collaboration et al., 2020): $\Omega_{m}=0.3111$, $\Omega_{\Lambda}=0.6889$, $\Omega_{b}=0.049$, $\sigma_{8}=0.8102$, and $n_{s}=0.9665$.
Dark matter haloes are identified using the friends-of-friends (FOF) algorithm (Davis et al., 1985) with a linking length of 0.2 times the mean interparticle separation. Subhaloes and their evolutionary histories are further processed using the Hierarchical Bound-Tracing (HBT+) algorithm (Han et al., 2012, 2018), which operates in the time domain to track the evolution of each halo throughout the simulation. The minimum number of particles in a subhalo is set to 20, corresponding to a minimum halo mass of approximately $7.5\times 10^{9}h^{-1}M_{\odot}$. The HBT+ code provides high-quality subhalo catalogues and a robust merger tree by following the most-bound particle even if a subhalo is no longer resolved.
We populate the dark matter halos in our simulation with mock galaxies using the Subhalo Abundance Matching (SHAM) method. This technique establishes a connection between the properties of galaxies and their corresponding subhaloes by assuming a monotonic relationship between these properties (Kravtsov et al., 2004; Vale & Ostriker, 2004; Conroy et al., 2006; Yang et al., 2012; Wechsler & Tinker, 2018). The positions and velocities of galaxies are assigned by directly inheriting the positions and velocities of their host subhaloes from the N-body simulations. The luminosities are simulated by matching the $z$-band cumulative galaxy luminosity functions measured from the DESI One-percent survey (DESI Collaboration et al., 2024a) with the cumulative subhalo mass functions derived from the peak mass of all the subhaloes. To reflect the luminosity-subhalo mass relation more realistically, a scatter in the $z$-band luminosity, $\sigma_{\mathrm{log}}(L_{z})=0.15$ dex (e.g. Yang et al., 2008), is added to each galaxy.
Figure 1: The footprint of DESI Legacy Imaging Survey DR9. The red contour indicates the region selected for reconstruction. The dashed black line represents the Galactic plane, excluding galaxies with $|b|<25$ in Galactic coordinates.
Figure 2: Normalized redshift distribution. Blue solid and red dashed lines correspond to the DESI Y1 catalog and our mock DESI sample, respectively.
We construct mock DESI survey and its corresponding dark matter particle distribution using the following steps:
  1. 1.
    We place a virtual observer at the center of the populated simulation box and stack replicas of the box to generate a past light cone. This process involves selecting snapshots from the N-body simulations that intersect with the light-cone volume at different redshifts.
  2. 2.
    We define an equatorial coordinate system and remove all mock galaxies or dark matter particles located outside the DESI LS DR9 footprint, as shown in Figure 1. We also apply a foreground mask to exclude regions containing globular clusters, planetary nebulae, nearby large galaxies, and Gaia stars with $G<16$. Additionally, we remove the region near the Galactic plane, $|b|<25$, to avoid areas of high stellar density.
  3. 3.
    For galaxies, we assign each a redshift, $z_{\rm obs}$, including the effects of redshift-space distortion, calculated as:
    $z_{\rm obs}=z_{\rm cos}+z_{\rm pec}=z_{\rm cos}+\frac{v_{\rm pec}}{c}(1+z_{\rm cos}),$ (1)
    where $z_{\rm cos}$ is the cosmological redshift from Hubble expansion; $z_{\rm pec}$ is the Doppler contribution due to the galaxy’s line-of-sight peculiar velocity; $v_{\rm pec}$, $c$ is the speed of light. For dark matter particles, we retain them in real space.
  4. 4.
    For galaxies, we assign a $z$-band apparent magnitude, $m_{z}$, based on their distance and luminosity, incorporating a $K$-correction described as:
    $K^{0.5}(z)=0.73z^{2}-0.54z-0.33.$ (2)
In this study, we focus on the BGS sample with a $z$-band magnitude limit cut $m_{z}<19.0$ and within the redshift range of $0.40.1<z<$ 0.4, which is consistent with the DESI DR1 BGS (DESI Collaboration et al., 2024b). Figure 2 shows the normalized redshift distributions for our mock DESI sample (blue solid curve). For comparison, we also include the redshift distribution derived from the DESI Y1 catalogue (black dashed line). The redshift distribution of our mock sample generally follows the observed trend. Since the DESI Y1 catalogue covers a relatively small fraction of the sky, a slight difference between the two distributions is expected due to cosmic variance. In addition, DESI will cover two sky regions: a larger region in the Northern Galactic Cap (NGC) and a smaller region in the Southern Galactic Cap (SGC). To use a more contiguous region, we apply our reconstruction on the mainland of NGC, as shown in the red contour of Figure 1. The final sky coverage used in our reconstruction is about 9625.23 deg2. In this study, we construct a set of 12 mock samples by randomly rotating and shifting the stacked simulation boxes, which are used for training and testing our network.
Figure 3: The projected field distributions in a slice of $1564\times 1564\times 9.16\,h^{-1}\,\mathrm{Mpc}$. From left to right, the panels correspond to the galaxy overdensity, dark matter density, and velocity magnitude $|\boldsymbol{v}|$, respectively.
To construct the grid-based fields for dark matter and galaxies within the DESI-like survey volume, we proceed as follows. First, we embed the survey volume within a periodic cubic box with a side length of $1564\,h^{-1}\,\mathrm{Mpc}$. To maximize the utilization of the box volume, we apply a constant angular rotation in the right ascension (R.A.) direction. The 3D positions $(x,y,z)$ of the objects are then calculated by
$x$ $=r\cos(\beta)\cos(\alpha-\alpha_{\mathrm{rot}}),$ (3)
$y$ $=r\cos(\beta)\sin(\alpha-\alpha_{\mathrm{rot}}),$
$z$ $=r\sin(\beta),$
where $r$ is the line-of-sight comoving distance, $\beta$ is the declination (Dec.), $\alpha$ is R.A., and $\alpha_{\mathrm{rot}}=58^{\circ}$ is the applied rotation angle in this study.
We use the cloud-in-cell (CIC) scheme (Hockney & Eastwood, 1981) to compute the overdensity field, $\delta$, on a grid of $512^{3}$ grids for both dark matter and galaxy distributions. The velocity field, $\boldsymbol{v}$, is also estimated using the CIC method and normalized by the number of dark matter particles in each cell of the grid. To suppress shot noise, we smooth both $\delta$ and $\boldsymbol{v}$ fields using a TensorFlow convolutional kernel with a kernel size of 3. Figure 3 shows the projected field distributions in a slice of $1564\times 1564\times 9.16\,h^{-1}\,\mathrm{Mpc}$. From left to right, the panels correspond to the galaxy overdensity, dark matter density, and velocity magnitude $|\boldsymbol{v}|$, respectively. The correlated structures among the three fields are clearly visible.
3 METHODOLOGY
In this section, we present our reconstruction method based on the UNet convolutional neural network.
3.1 Reconstruction via Maximum-Likelihood Estimation
Reconstructing the dark matter density and velocity fields from the galaxy density field constitutes a field-to-field mapping problem, which can be framed as a maximum likelihood estimation (MLE) task (e.g. Mao et al., 2021). The objective is to infer the underlying dark matter fields given the observed galaxy density field. Assuming a probabilistic relationship between the two, the likelihood function is given by,
$\mathcal{L}=\prod_{k=1}^{N_{p}}P(\phi_{\mathrm{m},k}|\delta_{\mathrm{g},k};\boldsymbol{\theta}),$ (4)
where $k$ indexes the grids, $N_{p}$ denotes the total number of grids, $\phi_{\mathrm{m}}$ represents the dark matter field (e.g., $\delta_{\mathrm{m}}$ or $\boldsymbol{v}_{\mathrm{m}}$), $\delta_{\mathrm{g}}$ is the galaxy density field, and $\boldsymbol{\theta}$ represents the model parameters. This formulation assumes that the grids are statistically independent, thereby simplifying the problem to a mapping between the corresponding grid points in the fields.
The estimation of $\boldsymbol{\theta}$ can be expressed in the following log-likelihood form,
[…]
where $f(\delta_{\mathrm{g},k};\boldsymbol{\theta})$ represents the predicted field value assuming a Gaussian distribution in the data. This shows that maximizing the likelihood is equivalent to minimizing the Mean Squared Error (MSE) between the true and predicted fields.
However, solving this formulation analytically is impractical due to the model’s complexity. Traditional optimization methods face challenges in estimating $\boldsymbol{\theta}$ due to the high dimensionality of the parameter space. Alternatively, neural networks provide an efficient approach by maximizing likelihood through supervised learning, effectively encoding $\boldsymbol{\theta}$ in their weights and biases.
Figure 4: Three-dimensional (3D) UNet architecture. Each cube represents a 3D space-with-feature map, with the grid size and the number of feature channels indicated. The input and output cubes, each containing $512^{3}$ voxels and spanning $1564\,\>h^{-1}{\rm{Mpc}}$, correspond to the redshift-space galaxy density field and the real-space dark matter density or velocity field, respectively. The encoder comprises 8 convolutional layers that progressively downsample the input cube to a bottleneck representation, while the decoder uses 8 transposed convolutional layers to upsample the bottleneck output to the target cube. The channel configuration (1+3) corresponds to the density field, $\delta_{\mathrm{m}}$, and the velocity components $(v_{x},v_{y},v_{z})$, enabling simultaneous reconstruction of these quantities as an optional feature. Computational steps are represented by distinct symbols, as illustrated in the lower-right corner. All convolutional layers utilize $4\times 4\times 4$ filters with a stride of $2$.
3.2 UNet Architecture
We base our network architecture on the U-Net style (Ronneberger et al., 2015), which is particularly well-suited for reconstructing complex fields due to its hierarchical convolutional layers and skip connections. These design features allow the model to capture multi-scale spatial information effectively, which is critical for learning the mapping between the redshift-space galaxy density field and the target fields.
To implement this architecture, we adapt the generator design from the Pix2Pix framework (Isola et al., 2016), tailoring it to our specific requirements. As illustrated in Figure 4, the network processes input and output cubes, each comprising $512^{3}$ grids. The input cube represents the redshift-space galaxy density field $\delta_{\mathrm{g}}$, while the output cube corresponds to either the real-space dark matter density field $\delta_{\mathrm{m}}$ or the velocity field $\boldsymbol{v}_{\mathrm{m}}$. The mapping is achieved through a symmetric encoder-decoder structure, where all convolutional layers utilize $4\times 4\times 4$ filters with a stride of $2$. The architecture includes the following key components:
  • Encoder Path: The encoder extracts features from the input field, $\delta_{\mathrm{g}}$, while progressively reducing spatial resolution. Each block consists of a sequence: convolution, batch normalization, and LeakyReLU activation. Batch normalization is omitted in the first and last layers of the encoder, where the last layer uses ReLU activation instead of LeakyReLU.
  • Decoder Path: The decoder reconstructs the target field incrementally by increasing spatial resolution. Each block consists of a sequence: transposed convolution, batch normalization, dropout (with a rate of 50%), and ReLU activation. Batch normalization is not applied in the first layer of the decoder, and the output layer employs a linear activation function.
  • Skip Connections: Skip connections link encoder layers to their corresponding decoder layers, enabling the preservation of spatial details during downsampling and upsampling.
  • Output Layer: For density field reconstruction, the output layer generates a single-channel output corresponding to $\delta_{\mathrm{m}}$. For velocity field reconstruction, the output layer produces three channels, representing the velocity components $(v_{x},v_{y},v_{z})$.
In this architecture, the fields $\delta_{\mathrm{m}}$ and $\boldsymbol{v}_{\mathrm{m}}$ are reconstructed separately due to memory limitations, while they could be reconstructed simultaneously by designing a multichannel output for the network if hardware permits.
Figure 5: (Upper left) Loss function for training $(\delta_{\mathrm{g}},\delta_{\mathrm{m}})$ using 3000 epochs. The dark solid and blue dashed lines represent the loss functions from the training and validation samples, respectively. (Upper right) Comparison of the density distributions for the validation samples. The black solid line represent the result of the target validation samples. Different colored lines represent the predicted fields at epochs 200, 1000, 2000, and 3000, repectively. (Lower left) Grid-to-grid density relation between the reconstructed density field $1+\delta_{\mathrm{rec}}$ and the true field $1+\delta_{\mathrm{true}}$. The contour in each case represents threshold encompassing 99% of the grid cells. The dashed lines indicate the ideal one-to-one correspondence. Different colored curves correspond to different epochs, as indicated. (Lower right) The evolution of the factor $f_{sc}$, defined as the product of the scatter variance and one minus the correlation scope between $\delta_{\mathrm{rec}}$ and $\delta_{\mathrm{true}}$, as a function of epoch.
3.3 Training Process
As described in Section 2, we generate paired data samples $(\delta_{\mathrm{g}},\delta_{\mathrm{m}})$ and $(\delta_{\mathrm{g}},\boldsymbol{v}_{\mathrm{m}})$ to train the networks. These samples represent the galaxy density field $\delta_{\mathrm{g}}$ paired with the true dark matter density field $\delta_{\mathrm{m}}$ or the velocity field $\boldsymbol{v}_{\mathrm{m}}$. The objective is to learn the mapping between the galaxy and dark matter fields.
The training process then minimizes the MSE loss function, which corresponds to the log-likelihood function in Equation (5). For practical implementation, the loss function is expressed as,
$\mathcal{L^{\prime}}=\frac{1}{N_{p}}\sum_{k=1}^{N_{p}}\left(\phi^{\mathrm{pred}}_{\mathrm{m},k}-\phi^{\mathrm{true}}_{\mathrm{m},k}\right)^{2},$ (6)
where $\phi^{\mathrm{pred}}_{\mathrm{m}}$ is the field being reconstructed and $\phi^{\mathrm{true}}_{\mathrm{m}}$ denote the true target field. Specifically, $\phi_{\mathrm{m}}\equiv\delta_{\mathrm{m}}$ for density reconstruction and $\phi_{\mathrm{m}}\equiv\boldsymbol{v}_{\mathrm{m}}$ for velocity reconstruction.
To improve the training stability and optimize the network’s ability to capture variations across a wide range of values, we rescale the input fields before training. The overdensity field is rescaled using a logarithmic transformation,
$y=\log_{10}(1+\delta+a),$ (7)
where the parameter $a=0.02$ ensures numerical stability by preventing singularities for low-density regions ($\delta\approx-1$). Moreover, the $a$ value is optimized in our analysis to balance the reconstruction, enhancing accuracy in high-density regions while preserving reliable performance in low-density areas. The velocity field is rescaled linearly as,
$y=\frac{\boldsymbol{v}_{\mathrm{m}}}{b},$ (8)
where the parameter $b=3000\,\mathrm{km}\,\mathrm{s}^{-1}$ normalizes the velocity values to a comparable scale, avoiding excessively large gradients during training. These rescaling transformations enhance the network’s convergence and improve its performance by standardizing the input dynamic range.
The network is trained from scratch using the Adam optimizer (Loshchilov & Hutter, 2017), which incorporates the adaptive learning rate. Each epoch processes 5 training samples, followed by 2 validation samples used exclusively for model evaluation during training.
In training neural networks, selecting an optimal stopping point is critical to balancing generalization and overfitting, particularly for fields with wide-range density distributions. As shown in the upper-left panel of Figure 5, the validation loss converges around epoch 200, indicating an appropriate point to terminate training. However, relying solely on loss metrics may overlook model performance across varying density regions, especially in high-density areas, which represent a small fraction of the total volume.
To address this limitation, we compare the density distributions of the predicted and true fields (upper-right panel of Figure 5). From epoch 200 to 2000, the disparities in high-density regions with $(1+\delta)>6$ progressively diminish, indicating significant improvements in recovering high-density structures even after loss convergence. However, as training progresses, predictions exhibit systematic underestimation at epoch 3000, revealing overfitting effects.
However, the grid-to-grid relationship (lower-left panel of Figure 5) support the opposite trend. From epoch 200 to 2000, the scatter increases while maintaining minimal bias, indicating improved density distribution recovery at the cost of reduced correlation with the true field. This trade-off highlights the need for a quantitative criterion to identify the optimal training stage.
We define the factor $f_{sc}$ as the product of the perpendicular variance, represented by the smaller eigenvalue of the covariance matrix of $(\vec{\delta}_{\mathrm{rec}},\vec{\delta}_{\mathrm{true}})$, and the absolute value of one minus the correlation slope, which is given by the slope of the eigenvector corresponding to the larger variance. As illustrated in the lower-right panel of Figure 5, this factor reaches its minimum value at approximately epoch 1300 and increases significantly after epoch 2000. Consequently, halting the training process around epoch 1300 achieves a balance between high-density recovery and correlation retention, ensuring the reconstructed field accurately represents cosmological structures.
These results underscore the necessity of carefully assessing the training stages to ensure the reliability of the model for cosmological inference. The methodology introduced here could be generalized to other datasets, offering a pathway to robust machine learning applications in astrophysical studies.
Figure 6: Comparison of the projected density-field distributions in a slice of $1564\times 1564\times 9.16\,h^{-1}\,\mathrm{Mpc}$. Different panels correspond to the redshift-space galaxy distribution (input), the true dark matter density field $\delta_{\mathrm{true}}$ (target), the UNet-reconstructed density field $\delta_{\mathrm{rec}}$ (output), and the difference between $\delta_{\mathrm{rec}}$ and $\delta_{\mathrm{true}}$, as labeled at the top of each panel.
Figure 7: (Left panel) Grid-to-grid density relationship between the reconstructed density field $1+\delta_{\mathrm{rec}}$ and the true field $1+\delta_{\mathrm{true}}$, represented by the blue solid line. For comparison, the result for the galaxy density field is also plotted with the red dashed line. The three contours in each case represent thresholds encompassing 67%, 95%, and 99% of the grid cells. The dashed lines indicate the ideal one-to-one correspondence. (Right panel) Density distributions for galaxies (red dashed), the reconstructed field (blue solid), and the true field (black solid). Error bars indicate the ±$\pm 1\sigma$ variance among the 5 test samples.
Figure 8: Same as Figure 7 but for comparing the densities in low-redshift (orange) and high-redshift region (blue).
Figure 9: Comparison of $P(k)$ (upper panel) and $P(k)$ ratios (lower panel). The black solid line and blue dashed line represent the auto-power spectrum of the true and reconstructed density fields respectively, while the green dot-dashed line represents their cross-power spectra. For reference, the red dotted line shows the auto-power spectra of the redshift-space galaxy field. The bottom panel shows the ratio between the reconstructed and true power spectra, with error bars indicating the ±$\pm\sigma$ variance across the five test samples. The black solid error bars are computed using Eq. (9).
Figure 10: The ratios $R_{\ell}(k)$ for the quadrupole (blue solid) and hexadecapole (orange dashed) power spectrum.
4 Results
In this section, we present the reconstructed density and velocity fields derived from the mock DESI survey by applying the resulting UNet model to 5 test samples. We also assess the reconstructed density field by validating its tidal field.
4.1 Reconstructing density field
For visual comparison, Figure 6 shows the projected density-field distribution in a slice of $1564\times 1564\times 9.16\,h^{-1}\,\mathrm{Mpc}$. Different panels correspond to the redshift-space galaxy distribution (upper left), the true dark matter density field $\delta_{\mathrm{true}}$ (upper right), the UNet-reconstructed density field $\delta_{\mathrm{rec}}$ (lower left), and the difference between $\delta_{\mathrm{rec}}$ and $\delta_{\mathrm{true}}$ (lower right). Compared with the true field, the reconstructed density field reveals recognizable and consistent large-scale structures, including clusters, filaments, and voids, indicating a successful reconstruction overall. Additionally, the results demonstrate resilience to boundary effects, as structures near the survey edges are effectively recovered. However, the difference map highlights discrepancies in high-density regions, particularly around clusters, where the reconstructed field deviates from the target.
The left panel of Figure 7 illustrates the grid-to-grid density relationship between the reconstructed density field $1+\delta_{\mathrm{rec}}$ and the true field $1+\delta_{\mathrm{true}}$, represented by the blue solid line. For comparison, the result for the galaxy density field is also plotted. The three contours in each case represent thresholds encompassing 67%, 95%, and 99% of the grid cells. The dashed lines indicate the ideal one-to-one correspondence. As expected, galaxies trace the underlying dark matter with a significant bias, leading to a relationship that deviates substantially from unity. However, after UNet reconstruction, the relationship becomes nearly linear and does not show significant bias.
The right panel of Figure 7 shows the density distributions for galaxies (red dashed), the reconstructed field (blue solid), and the true field (black solid). The error bars indicate the ±$\pm 1\sigma$ variance among the 5 test samples. After reconstruction, $\delta_{\mathrm{rec}}$ closely matches $\delta_{\mathrm{true}}$ in the regions where $71+\delta<$ 7, but a systematic underestimation occurs for $71+\delta>$ 7. This discrepancy likely arises from the imbalance in the number of grid cells between low- and high-density regions. As discussed in Section 3.3, the network initially prioritizes accurate predictions in low-density regions during early training stages, gradually improving high-density recovery at the cost of reduced correlation with the true field in later stages. The final model was selected by balancing scatter and correlation measures, which may contribute to the observed divergence. However, this issue affects only about 0.2% of the grids with the $71+\delta>$ 7 region. In comparison, the galaxy density distribution fails to trace the dark matter field accurately across both low- and high-density regimes.
Additionally, Figure 8 presents the comparison between the low-redshift ($0.1\lesssim z\lesssim 0.2$) and high-redshift ($0.3\lesssim z\lesssim 0.4$) regions. In the left panel, the high-redshift region exhibits slightly larger scatter, highlighting the impact of lower galaxy density. In the right panel, the reconstructed density distribution underestimates high-density regions more significantly in the low-redshift region. This underestimation likely arises from higher galaxy clustering at low redshifts, which enhances nonlinear effects that the reconstruction method cannot fully correct, resulting in a suppression of high-density peaks. Overall, the method demonstrates robust performance, though further refinement is needed to better address nonlinear effects.
To evaluate the clustering, we compare the power spectrum of the reconstructed dark matter density fields with the true fields. In this study, random samples were not used to correct for the survey’s geometry or selection effects, as the focus is on evaluating the reconstruction’s fidelity through direct comparison. This highlights the quality of the reconstruction but does not represent the true underlying clustering.
The top panel of Figure 9 illustrates the power spectrum comparisons. The black solid line and blue dashed line represent the auto-power spectra of the true and reconstructed density fields, respectively. For reference, the red dotted line shows the auto-power spectrum of the redshift-space galaxy density field. The bottom panel shows the ratio between the reconstructed and true power spectra, with error bars indicating the ±$\pm\sigma$ variance across the five test samples. By considering the cosmic variance due to the limited number of Fourier modes for a finite-volume survey, we presented the theoretical error (black error bar) for the power spectrum(e.g. Song et al., 2021) by,
$\Delta P(k)=\frac{P(k)}{\sqrt{N_{k}}}$ (9)
where $N_{k}$ denotes the number of independent $k$-modes available per bin. Compared to the halo $P(k)$, the reconstructed autocorrelation $P(k)$ values align with the true power spectrum within $1\sigma$ across the range of scales $k<0.1\,h\,\mathrm{Mpc}^{-1}$. The underestimation on scale $k>0.1~{}h~{}\mathrm{Mpc}^{-1}$ is probably due to inaccurate reconstruction in the high-density region, as discussed above. The model’s training objective, typically centered on minimizing global error metrics, may prioritize large-scale accuracy over small-scale fidelity, causing a modest degradation in the reconstructed small-scale power.
Figure 11: Comparison between the true and reconstructed dark-matter velocity fields in a slice of dimensions $1564\times 1564\times 3.05\,h^{-1}\,\mathrm{Mpc}$. The left panel shows the true velocity field $\vec{v}_{\mathrm{true}}$, directly extracted from the original simulation, while the middle panel presents the UNet-reconstructed field $\vec{v}_{\mathrm{rec}}$. The right panel shows the residuals, $\vec{v}_{\mathrm{rec}}-\vec{v}_{\mathrm{true}}$. In all panels, the arrows represent the velocity vectors, with their lengths proportional to the vector magnitudes. The underlying density field is also presented.
To evaluate the concordance between the reconstructed and true density fields, we analyzed the cross-correlation power spectrum. Unlike the auto-power spectrum, which assesses the overall clustering amplitude, the cross-power spectrum directly measures how well the reconstructed field reproduces the phase and spatial distribution of the true field. High cross-correlation values indicate that the reconstruction effectively captures both large- and small-scale structures, making it a robust validation metric. In Figure 9, the green dash-dotted line represents the cross-power spectrum of the true and reconstructed fields, as well as the ratios of the reconstructed cross-power spectrum to the true one. The results show that the ratios align with unity within the $1\sigma$ range on scales $k<0.1\,h\,\mathrm{Mpc}^{-1}$, indicating a strong agreement between the two fields. At smaller scales ($k>0.1\,h\,\mathrm{Mpc}^{-1}$), the ratios fall below unity. Specifically, the reductions in the cross-correlation power spectrum are approximately 1.5% at $k=0.1\,h\,\mathrm{Mpc}^{-1}$ and 20% at $k=0.3\,h\,\mathrm{Mpc}^{-1}$. These reductions, though moderate, reflect challenges in accurately reconstructing high-density regions at smaller scales. Quantitatively, the average cross-correlation coefficient remains above 0.985 for $k<0.1\,h\,\mathrm{Mpc}^{-1}$, confirming the high fidelity of the reconstructed field for large-scale structures.
To evaluate the accuracy of redshift-space distortion (RSD) corrections in the reconstructed dark matter density field, we compute the ratios of the multipole moments of the power spectra. These ratios compare the reconstructed and true power spectra, normalized by the difference between the redshift-space galaxy density power spectrum and the true power spectrum. The general formula for the ratio for any multipole $\ell$ is expressed as:
$R_{\ell}(k)=\frac{\lvert P_{\ell}^{\text{rec}}(k)-P_{\ell}^{\text{true}}(k)\rvert}{\lvert P_{\ell}^{\text{gal}}(k)-P_{\ell}^{\text{true}}(k)\rvert},$ (10)
where $P_{\ell}^{\text{rec}}(k)$ is the reconstructed dark-matter power spectrum, $P_{\ell}^{\text{true}}(k)$ is the true dark-matter power spectrum, and $P_{\ell}^{\text{gal}}(k)$ is the redshift-space galaxy density power spectrum. The ratios are computed for both the quadrupole ($\ell=2$) and hexadecapole ($\ell=4$), providing a measure of the RSD correction performance at different scales.
Figure 10 shows the ratios $R_{\ell}(k)$ for both the quadrupole ($\ell=2$) and hexadecapole ($\ell=4$) power spectrum. At large scales, the ratios are close to zero, indicating that the RSD correction effectively restores the reconstructed power spectra to the true values. At small scales, the ratios are greater than zero, but typically remain less than 0.5. This suggests that most of the RSDs have been successfully removed, although some residual distortions persist at smaller scales, probably due to less effective reconstruction in high-density regions.
Based on the test results, we conclude that our network, despite using a small sample size, is capable of directly predicting the density field from the galaxy density field and effectively correcting for redshift-space distortions (RSDs). This leads to a reliable reconstruction of the clustering dark-matter density field. However, it is important to note that additional training data is still needed to enhance performance, particularly in high-density regions and on smaller scales.
Figure 12: (Upper panels) Grid-to-grid velocity relations for the $x$-, $y$-, and $z$-axis velocity components, as shown in different columns. The blue solid line corresponds to the results from the entire survey region, while the red dashed lines represent the results for the center survey region, defined as a sub-box within $391.0,h^{-1},\mathrm{Mpc}<x,y,z<1173.0,h^{-1},\mathrm{Mpc}$. The contours represent 67%, 95%, and 99% of the grid cells, with the best-fitting linear regressions shown as dashed lines. The slope of the best-fitting relation and the scatter are also indicated in each panel. (Lower panels) Probability distributions of velocity differences, $v_{\mathrm{rec}}-v_{\mathrm{true}}$.
Figure 13: Comparison of the velocity-field power spectrum. From left to right, the panels correspond to the velocity magnitude, divergence, and the vorticity, respectively. In each panel, the black solid and blue dashed lines represent the power spectrum of the true and reconstructed velocity field, respectively. The green dot-dashed line shows the cross-correlation power spectra for the velocity magnitude between the true and reconstructed fields.
4.2 Reconstructing the velocity field
Figure 11 illustrates the comparison between the true and reconstructed dark-matter velocity fields in a slice of dimensions $1564\times 1564\times 3.05\,h^{-1}\,\mathrm{Mpc}$. The left panel shows the true velocity field, $\vec{v}_{\mathrm{true}}$, directly extracted from the original simulation, while the middle panel presents the UNet-reconstructed field, $\vec{v}_{\mathrm{rec}}$. The right panel shows the residuals, $\vec{v}_{\mathrm{rec}}-\vec{v}_{\mathrm{true}}$. In all panels, the arrows represent the velocity vectors, with their lengths proportional to the vector magnitudes. The underlying density field is also presented. The reconstructed field $\vec{v}_{\mathrm{rec}}$ closely matches $\vec{v}_{\mathrm{true}}$, successfully capturing both large-scale coherent flows and small-scale turbulent features. This demonstrates the UNet’s capability to recover the true velocity structure across a wide range of scales. However, the residual map on the right panel highlights discrepancies mainly near the survey boundaries. These deviations are likely caused by boundary effects, as the reconstruction model struggles to fully account for the lack of surrounding information at the edges of the domain. Despite these boundary-related residuals, the overall performance of the reconstruction remains robust, with the primary flow patterns and magnitudes well-reproduced.
The upper panels of Figure 12 show the grid-to-grid velocity relations for the $x$-, $y$-, and $z$-axis velocity components, as shown in different columns. The blue solid line corresponds to the results from the entire survey region, while the red dashed lines represent the results for the center survey region, defined as a sub-box within $391.0,h^{-1},\mathrm{Mpc}<x,y,z<1173.0,h^{-1},\mathrm{Mpc}$. The contours represent 67%, 95%, and 99% of the grid cells, with the best-fitting linear regressions shown as dashed lines. The slopes of the best-fit relations are close to unity for both regions, with values around 0.950.950.950.951.001.001.001.00, indicating strong agreement between the reconstructed $v_{\mathrm{rec}}$ and true velocities $v_{\mathrm{true}}$. The scatter values, reflecting the variances from the best-fit lines, are smaller in the center region, demonstrating improved reconstruction precision due to reduced boundary effects. The lower panels of Figure 12 illustrate the probability distributions of velocity differences, $v_{\mathrm{rec}}-v_{\mathrm{true}}$. These distributions peak nicely at zero, and the center region shows narrower peaks, further verifying its superior accuracy. In contrast, the broader distributions in the entire survey region underscore the challenges posed by boundary effects and edge distortions. Overall, the results demonstrate the reliability of the reconstruction, with notable improvements in the center survey region compared to the full survey.
To further evaluate the reconstructed velocity field, we calculate the power spectrum, as shown in Figure 13. In each panel, the black solid line represents the power spectrum of the true velocity field, while the blue dashed line corresponds to the reconstructed field. Additionally, the green dot-dashed line shows the cross-correlation power spectra for the velocity magnitude between the true and reconstructed fields, which serves as a measure of their concordance. Overall, the reconstructed velocity field demonstrates strong agreement with the true field across the three components, with some deviations observed.
The cross-power spectrum for the velocity magnitude is systematically lower than unity by approximately 10% on all scales. This discrepancy can be attributed to two factors: (1) survey boundary effects, which truncate the velocity field at the edges of the survey volume and introduce artifacts, particularly on large scales, weakening the correlation between the true and reconstructed modes; and (2) training limitations, as the reconstruction model, typically trained to minimize global error metrics, prioritizes large-scale accuracy at the cost of small-scale fidelity, resulting in suppressed power on smaller scales. For the vorticity power spectrum, the discrepancy at scales $k<0.2\,h\,\mathrm{Mpc}^{-1}$ may arise from the intrinsical nature of the vorticity signal. As vorticity is primarily sourced from small-scale, nonlinear velocity interactions, the reconstruction method may face challenges in accurately reproducing these subtle features at relatively large scale where the signal-to-noise ratio is inherently low. However, at smaller scales $k>0.2\,h\,\mathrm{Mpc}^{-1}$, the reconstructed vorticity shows better agreement with the true field, reflecting the model’s ability to capture localized turbulence. These results underscore the robust overall performance of the reconstruction method, while also highlighting the limitations introduced by survey boundaries and algorithmic challenges.
Figure 14: Eigenvalues of the reconstructed tidal tensor versus those obtained from the true dark-matter density field. The contours for each result encompass 95% and 99% of the grid cells.
Figure 15: Comparison of the classification of the large-scale structure between the true dark-matter density field (left panel) and UNet-reconstructed density field (right panel) in a slice of $1564\times 1564\times 3.05\>h^{-1}{\rm{Mpc}}$. The yellow, yellow-green, green, and black regions are grid cells located at structures classified as clusters, filaments, sheets, and voids, respectively.
4.3 Reconstructing the tidal field
We now proceed with the reconstruction of the large-scale tidal field using the tidal tensor, $T_{ij}$, defined as
$T_{ij}=\frac{\partial^{2}\phi}{\partial x_{i}\,\partial x_{j}},$ (11)
where $i$ and $j$ are indices taking values 1, 2, or 3, and $\phi$ is the peculiar gravitational potential. We computed $\phi$ from the mass density field using the Poisson equation,
$\nabla^{2}\phi=4\pi G\bar{\rho}\delta,$ (12)
where $\bar{\rho}$ is the average density of the universe. The tidal tensor is then diagonalized to obtain the eigenvalues $\lambda_{1}\geqslant\lambda_{2}\geqslant\lambda_{3}$ at each grid point.
Figure 14 compares the eigenvalues of the tidal field tensor ($T_{1}$, $T_{2}$, $T_{3}$) derived from the UNet-reconstructed dark-matter density field to those obtained from the true dark-matter density field of the mock DESI survey. The relation plots for each eigenvalue reveal a strong alignment along the diagonal, demonstrating excellent agreement between the reconstructed and true tidal fields. The contours, representing 95% and 99% of the grid cells, further confirm that the reconstructed eigenvalues closely match the true values, validating the accuracy of the reconstruction method. This result is consistent with our previous work (Wang et al., 2024), which showed that the UNet-based reconstruction method significantly outperforms traditional halo-based approaches (Wang et al., 2009) in recovering unbiased tidal fields.
The eigenvalues of the tidal field can be used to classify the large-scale structure (LSS) into four categories based on their local morphology:
  • Cluster: grid cells with three positive eigenvalues;
  • Filament: grid cells with one negative and two positive eigenvalues;
  • Sheet: grid cells with two negative and one positive eigenvalue;
  • Void: grid cells with three negative eigenvalues.
Figure 15 shows the LSS classifications derived from the true dark-matter density field (left panel) and the UNet-reconstructed density field (right panel). The visual comparison highlights a remarkable consistency between the two panels, with the spatial distributions of structures, including clusters, filaments, sheets and voids, closely matching across the slice. This indicates that the UNet-reconstructed dark-matter density field effectively preserves the topological connectivity and morphological details of the cosmic web. However, small discrepancies are observed, including a slight underestimation of the void volume near the boundaries and misrepresentation of filamentary or sheet regions. These differences likely arise from boundary effects in the survey region, which can truncate the density field near the edges, as well as the inherent challenges in predicting intermediate-density environments, where the contrast between sheets and filaments is less distinct. Despite these minor biases, the UNet reconstruction provides a robust and reliable representation of the tidal field, enabling accurate LSS classification for cosmological studies.
5 Summary and Conclusion
In this study, we reconstructed the dark matter density, velocity, and tidal fields using a DESI-like Bright Galaxy Sample (BGS) based on a UNet convolutional neural network. Our approach transforms galaxy density fields into the underlying dark matter properties by exploiting the U-Net’s capability to establish precise field-to-field mappings. This method significantly enhances the recovery of large-scale structures and captures the intricate morphology of cosmic web features.
We first generated a DESI-like survey using the Jiutian N-body simulation, which provides high resolution in accordance with Planck2018 cosmological parameters. From the survey, we selected BGS galaxies and the underlying dark matter particles within a redshift range of $0.40.1<z<$ 0.4. The resulting lightcone was carefully trimmed to DESI’s sky coverage of approximately $9625.23\,\mathrm{deg^{2}}$, providing a realistic and comprehensive dataset for both galaxy and dark matter distributions to train and test our models.
We successfully reconstructed the dark matter density field, achieving close agreement with the true field across large-scale structures such as clusters, filaments, and voids. The cross-correlation power spectrum demonstrates a reconstruction accuracy exceeding 98.5% for scales with $k<0.1\,h\,\mathrm{Mpc}^{-1}$. Moderate discrepancies, with reductions of 1.5% at $k=0.1\,h\,\mathrm{Mpc}^{-1}$ and 20% at $k=0.3\,h\,\mathrm{Mpc}^{-1}$, were observed in high-density regions due to challenges in reconstructing small-scale features. Furthermore, we evaluated the performance of the RSD correction using multipole moment ratios. These results confirm that the reconstructed field effectively restores the clustering signal by removing most of the RSD effects, achieving strong consistency with the true field, particularly at large scales.
Our velocity field reconstruction accurately recovered large-scale coherent flows and small-scale turbulent features. Grid-to-grid comparisons yielded slopes close to unity (0.95–1.00) with reduced scatter in central regions, highlighting robust performance. The cross-correlation power spectrum was systematically underestimated by approximately 10% across all scales, primarily due to boundary effects and small-scale reconstruction challenges. Despite these limitations, the results emphasize the model’s ability to retain the structure of the velocity field on a wide range of scales.
We reconstructed the tidal field by computing the eigenvalues of the tidal tensor and found strong agreement between the reconstructed and true fields. Our approach accurately preserved large-scale structure classifications, including clusters, filaments, sheets, and voids, while maintaining the morphological connectivity of the cosmic web. Minor biases were observed, such as a slight underestimate of the void volume near the boundaries and a misrepresentation of filamentary or sheet regions. These discrepancies are mainly attributed to boundary effects and the challenges of predicting intermediate-density environments. In general, the results confirm the ability of the method to reliably recover the structure of the tidal field.
We have demonstrated that deep learning, specifically a UNet-based architecture, is an effective tool for reconstructing cosmological fields from galaxy samples. Our method consistently achieved high fidelity across density, velocity, and tidal fields, providing a reliable framework for large-scale structure analysis. This study highlights potential applications in observational surveys and enhances our understanding of dark matter distributions and dynamics.
This work is supported by the National SKA Program of China (2022SKA0110200 and 2022SKA0110202), the National Key R&D Program of China (2023YFA1607800, 2023YFA1607801, 2023YFA1607804), the National Natural Science Foundation of China (Nos.12103037, 12261141691, 12373005), 111 project No. B20019, Shanghai Natural Science Foundation (Nos. 19ZR1466800, 24DX1400100, 23JC1410200) and Zhangjiang special support (No. ZJ2023-ZD-003). We acknowledge the science research grants from the China Manned Space Project (Grant No. CMS-CSST-2021-A02, CMS-CSST-2021-A03, CMS-CSST-2021-B01) and the Fundamental Research Funds for the Central Universities (Grant No. XJS221312). H.W. acknowledgements supports from the National Natural Science Foundation of China (NSFC, Nos. 12192224) and CAS Project for Young Scientists in Basic Research, Grant No. YSBR-062. The authors acknowledge the Beijing Super Cloud Center (BSCC) for providing HPC resources that have contributed to the research results reported in this paper. This work is also supported by the High-Performance Computing Platform of Xidian University.